OpenStudy (anonymous):

divide 9x/2a divide 3x^4/4ax simplify as much as possible

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \cfrac{9x}{2a}\div \cfrac{3x^4}{{\color{blue}{ 4ax}}}\implies \cfrac{9x}{2a}\cdot \cfrac{{\color{blue}{ 4ax}}}{3x^4}\implies \cfrac{3\cdot 3\cdot x\cdot 4\cdot a\cdot x}{2\cdot a\cdot 3\cdot x\cdot x\cdot x\cdot x}\) simplify away

OpenStudy (jdoe0001):

well...

OpenStudy (jdoe0001):

\(\bf \cfrac{9x}{2a}\div \cfrac{3x^4}{{\color{blue}{ 4ax}}}\implies \cfrac{9x}{2a}\cdot \cfrac{{\color{blue}{ 4ax}}}{3x^4}\implies \cfrac{3\cdot 3\cdot x\cdot 2\cdot 2\cdot a\cdot x}{2\cdot a\cdot 3\cdot x\cdot x\cdot x\cdot x}\) there, simplify away

OpenStudy (anonymous):

6x^2a/ax^2?

OpenStudy (jdoe0001):

well... \(\bf \cfrac{9x}{2a}\div \cfrac{3x^4}{{\color{blue}{ 4ax}}}\implies \cfrac{9x}{2a}\cdot \cfrac{{\color{blue}{ 4ax}}}{3x^4}\implies \cfrac{3\cdot \cancel{3}\cdot \cancel{x}\cdot \cancel{2}\cdot 2\cdot \cancel{a}\cdot \cancel{x}}{\cancel{2}\cdot \cancel{a}\cdot \cancel{3}\cdot x\cdot x\cdot \cancel{x}\cdot \cancel{x}}\)

OpenStudy (anonymous):

6/x^2

OpenStudy (jdoe0001):

yeap